A father is 5 times as old as his son. In 10 years he will be three times as old as his son. Form an equation and solve it to find his father's age

Answer by Cromlix(4381)

A father is 5 times as old as his son.
F = 5S
In 10 years he will be three times as old as his son.
F + 10 = 3(S + 10)
replace F with %S from top equation.
5S + 10 = 3(S + 10)
5S + 10 = 3S + 30
Collect like terms
5S - 3S = 30 - 10
2S = 20
S = 10
Currently;
Son is 10 years old
Father is 50 years old.
Hope this helps.
:-)